### v3

# Orthogonalized Policy Optimization Decoupling Sampling Geometry from Optimization Geometry in RLHF

Abstract
Large Language Model (LLM) alignment objectives are often presented as a collection of distinct algorithms—PPO, DPO, IPO, and their variants—each motivated by different derivations. In this work, we argue that this diversity obscures a simpler underlying structure. At a fundamental level, alignment objectives involve two independent design choices: (i) how training signals are sampled and weighted, and (ii) how deviations from a reference policy are geometrically penalized. Existing methods typically entangle these choices through a single divergence, most commonly the Kullback–Leibler (KL) divergence.
We show that this entanglement is not merely a modeling convenience but a source of systematic instability. When the same divergence simultaneously determines sample weighting and optimization curvature, adjusting one aspect (e.g., exploration strength) inevitably alters the other (e.g., gradient geometry). This coupling is particularly problematic in preference-based reinforcement learning, where advantage signals are unbounded and high-confidence regimes are common.
We propose a principled structural remedy: formulate alignment as an orthogonal mirror descent problem, in which sampling geometry enters as a linear driving force derived from an $\alpha$-divergence projection, while optimization geometry is determined independently by a Bregman divergence (mirror map). This perspective leads to a new alignment objective—Orthogonalized Policy Optimization (OPO)—obtained by choosing a Euclidean mirror map in likelihood-ratio space. The resulting objective admits a closed-form solution, linear (non-saturating) gradient dynamics, and a well-conditioned trust region, while remaining fully compatible with standard LLM training pipelines.
1 Introduction
Aligning Large Language Models (LLMs) with human preferences has become a cornerstone of modern AI development. The standard paradigm, Reinforcement Learning from Human Feedback (RLHF) [5, 6], typically employs algorithms like Proximal Policy Optimization (PPO) [2] or Direct Preference Optimization (DPO) [1] to optimize a policy against a reward model or preference dataset. Despite their formulation differences, these methods share a common structural foundation: they all rely on Kullback–Leibler (KL) divergence to constrain the policy update, preventing it from deviating excessively from a reference distribution.
While empirically effective, KL-regularized objectives exhibit distinct limitations in reasoning-intensive domains where high confidence is required. As the policy improves and assigns high probability to correct reasoning chains, the KL penalty—which induces an exponential geometry in the log-probability space—often dominates the learning signal. This manifests as gradient saturation, where the driving force for further improvement vanishes exponentially as the model becomes confident. Consequently, training often plateaus prematurely, behavior commonly attributed to "over-regularization" but which we argue is intrinsic to the chosen geometry.
We contend that this issue stems from a structural conflation in existing objective designs. Fundamentally, alignment involves two independent design choices: sampling geometry, which determines the effective weighting of training examples (e.g., whether to focus on high-advantage samples), and optimization geometry, which determines the curvature of the update step (e.g., how to measure distance in the policy space). In standard KL-based methods, a single divergence term dictates both, coupling the exploration strength with the optimization stability. Adjusting one inevitably perturbs the other, creating a dilemma where aggressive sampling destabilizes training, while stable optimization stifles exploration.
To resolve this, we propose a unifying framework based on Orthogonal Mirror Descent. By viewing alignment as a mirror descent process, we can explicitly decouple these two axes: the sampling geometry acts as a linear driving force defined by an advantage-weighted distribution, while the optimization geometry is independently governed by a Bregman divergence (mirror map). This perspective reveals that the saturation observed in DPO and PPO is not a necessary feature of alignment, but a specific consequence of using the Negative Entropy mirror map (which induces KL divergence).
Guided by this framework, we introduce Orthogonalized Policy Optimization (OPO), a new alignment objective derived by selecting the Euclidean mirror map in the space of likelihood ratios. Unlike the exponential geometry of KL, the Euclidean geometry induces a Pearson $\chi^{2}$ trust region, leading to a quadratic regularizer that maintains a linear, non-saturating driving force even in high-confidence regimes. This formulation cleanly separates the sampling intensity (controlled by an $\alpha$-divergence parameter) from the regularization strength (controlled by a stiffness parameter), allowing for robust, continuous improvement.
Our contributions are as follows:
  • We identify the implicit coupling of sampling and optimization geometries in KL-based alignment methods as a root cause of gradient saturation.
  • We propose the Orthogonal Mirror Descent framework, which decouples these design axes via two independent geometries: an $\alpha$-divergence projection governs sampling, while a Bregman divergence (mirror map) governs optimization.
  • We derive Orthogonalized Policy Optimization (OPO), a simple, closed-form objective based on Euclidean geometry that ensures linear gradient dynamics and global contraction.
  • We empirically demonstrate that OPO outperforms strong baselines like DPO and GRPO on mathematical reasoning tasks, validating the benefits of non-saturating gradients.
2 Related Work
Preference Optimization and RLHF.
Reinforcement Learning from Human Feedback (RLHF) typically involves learning a reward model from preferences and then optimizing a policy via PPO [2, 5, 6]. Direct Preference Optimization (DPO) [1] simplifies this by deriving a closed-form solution to the KL-constrained reward maximization problem. Recent variants extend this paradigm: IPO [7] adds a regularization term to prevent overfitting, SimPO [8] simplifies the reference-free objective.
$f$-Divergences in Machine Learning.
[…]
Trust-Region Methods.
TRPO [3] enforces stability via explicit KL constraints, while PPO [2] approximates this with ratio clipping. ADPO [17] shows that anchored coordinates provide an implicit trust region via temperature-scaled curvature. OPO extends this by replacing the KL-based geometry with a quadratic ($\chi^{2}$) geometry in ratio coordinates, providing a different form of implicit regularization.
3 Preliminaries: Two Coordinate Systems
Before presenting the OPO framework, we establish the two coordinate systems that underlie all ratio-based alignment objectives.
3.1 Ratio Space vs. Log-Ratio Space
Let the reference policy be $\pi_{\mathrm{ref}}$. We define two coordinate systems for measuring policy deviation:
Definition 3.1 (Ratio Coordinates).
The density ratio and centered ratio are:
$t_{\theta}(y):=\frac{\pi_{\theta}(y)}{\pi_{\mathrm{ref}}(y)},\qquad v_{\theta}(y):=t_{\theta}(y)-1=\frac{\pi_{\theta}(y)}{\pi_{\mathrm{ref}}(y)}-1$ (1)
Definition 3.2 (Log-Ratio Coordinates).
The log-ratio (natural parameterization of LLMs) is:
$\Delta_{\theta}(y):=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$ (2)
so that $t_{\theta}(y)=\exp(\Delta_{\theta}(y))$ and $v_{\theta}(y)=\exp(\Delta_{\theta}(y))-1$.
The key observation is that $\chi^{2}$ divergence is naturally quadratic in ratio coordinates $v$, while KL divergence involves the log-ratio $\Delta$. This distinction is crucial for understanding the gradient dynamics of different alignment objectives.
4 Theoretical Framework: Orthogonal Mirror Descent
We present a unified framework that fundamentally decouples the preference sampling geometry (what to learn) from the optimization geometry (how to learn). We formulate policy alignment as a functional Orthogonal Mirror Descent problem.
4.1 The General Framework
We view alignment as an iterative projection process. At each step $k$, we seek a policy $\pi$ that maximizes the expected return under a specific sampling distribution $\rho_{\alpha}$, subject to a trust-region constraint defined by a divergence $D_{\Psi}$:
$\pi_{k+1}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{y\sim\rho_{\alpha}}\left[A(y)\log\pi(y)\right]-\frac{1}{\eta}D_{\Psi}(\pi\|\pi_{k})$ (3)
This formulation cleanly orthogonizes two design axes:
  1. 1.
    Sampling Geometry ($\rho_{\alpha}$): Determines the effective weighting of samples (the linear driving force).
  2. 2.
    Optimization Geometry ($D_{\Psi}$): Determines the mirror map (regularization/curvature) used to project the update.
4.2 Axis 1: Sampling Geometry via $\alpha$-Divergence Projection
The sampling weight $\omega_{\alpha}$ determines the effective weighting of training examples. We derive it from a principled $\alpha$-divergence interpolation, providing the sampling axis with the same geometric rigor as the optimization axis.
Definition 4.1 (Oracle Target Distribution).
Given a non-negative quality signal $Q(y)\geq 0$ and the behavior policy $\tilde{p}(y):=\pi_{\mathrm{old}}(y)$, define the oracle target distribution:
$P^{*}(y)\;\propto\;Q(y)$ (4)
In the standard RL setting with advantage function $A(y)$, the natural choice is $Q(y)=\exp(A(y))$, corresponding to the Boltzmann-optimal density. The importance sampling (IS) ratio is $w_{\mathrm{IS}}(y)=P^{*}(y)/\tilde{p}(y)\propto Q(y)/\tilde{p}(y)$.
Direct use of $w_{\mathrm{IS}}$ suffers from variance explosion in large action spaces. The parameter $\alpha\in[0,1]$ geometrically tempers this correction by interpolating along the $\alpha$-geodesic [11] connecting $\tilde{p}$ to $P^{*}$:
Proposition 4.2 ($\alpha$-Escort Weighting).
The $\alpha$-geodesic between $\tilde{p}$ and $P^{*}$ yields the escort distribution $\rho_{\alpha}(y)\propto\tilde{p}(y)^{\alpha}\,P^{*}(y)^{1-\alpha}$. The corresponding importance weight is:
$\boxed{\omega_{\alpha}(y)\;=\;\frac{\rho_{\alpha}(y)}{\tilde{p}(y)}\;\propto\;\left(\frac{P^{*}(y)}{\tilde{p}(y)}\right)^{1-\alpha}=\left(\frac{\exp(A(y))}{\tilde{p}(y)}\right)^{1-\alpha}}$ (5)
This parameterization has clear geometric semantics:
  • $\alpha\to 1$: $\omega_{\alpha}\to 1$ (uniform weighting; mode-covering, conservative).
  • $\alpha\to 0$: $\omega_{\alpha}\to P^{*}/\tilde{p}$ (full IS correction; peak-seeking, aggressive).
  • $0<\alpha<1$: Smooth interpolation governed by $\alpha$-geometry.
Remark 4.3 (On-Policy Implementation with Frozen Denominator).
In the on-policy setting where $\tilde{p}=\pi_{\mathrm{old}}$, the theoretical escort weight $\omega_{\alpha}=(\exp(A)/\tilde{p})^{1-\alpha}$ can be decomposed as:
$\omega_{\alpha}(y)\;=\;\exp\!\big((1-\alpha)\,A(y)\big)\;\cdot\;\operatorname{sg}\!\left(\frac{1}{\pi_{\mathrm{old}}(y)^{\,1-\alpha}}\right)$ (6)
where $\operatorname{sg}(\cdot)$ denotes the stop-gradient operator. This formulation is not an approximation—it computes the exact escort weight from Eq. (5). The stop-gradient on the $1/\pi_{\mathrm{old}}^{1-\alpha}$ factor serves two critical purposes: (i) it prevents the $1/p$ denominator from injecting unbounded gradients, preserving the constant Hessian $\nabla^{2}_{v}\mathcal{L}=\mu I$; and (ii) it is semantically natural, since $\pi_{\mathrm{old}}$ is already frozen as the reference policy in each iteration.
Remark 4.4 (Two-Geometry Structure).
With this derivation, the entire OPO framework admits a clean geometric decomposition:
Alignment == $\alpha$-geometry (sampling) $\times\;$ mirror geometry (optimization)
The sampling axis is an $\alpha$-divergence interpolation in the statistical manifold; the optimization axis is a Bregman divergence projection in ratio space. These two geometries are structurally orthogonal: $\alpha$ shapes the driving force without altering the Hessian, and $\mu$ shapes the curvature without altering the target.
4.3 Axis 2: Optimization Geometry: Euclidean Mirror Map
The core innovation of OPO is the choice of the mirror map. Instead of the standard Negative Entropy map (which leads to KL divergence and exponential updates), we choose the Euclidean Mirror Map in the space of likelihood ratios $v=\pi/\pi_{\mathrm{ref}}-1$:
$\Psi(v)=\frac{1}{2}\|v\|_{L^{2}(\pi_{\mathrm{ref}})}^{2}=\frac{1}{2}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}[v(y)^{2}]$ (7)
The induced Bregman divergence is the Pearson $\chi^{2}$ divergence:
$D_{\Psi}(\pi\|\pi_{k})=\frac{1}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\left[\left(\frac{\pi}{\pi_{\mathrm{ref}}}-\frac{\pi_{k}}{\pi_{\mathrm{ref}}}\right)^{2}\right]=D_{\chi^{2}}(\pi\|\pi_{k})$ (8)
This choice transforms the alignment problem into a quadratic optimization in ratio space, with curvature $\mu I$ independent of the data.
4.4 The OPO Objective
Combining the $\alpha$-escort sampling weight (Axis 1) with the Euclidean mirror map (Axis 2) yields the OPO objective:
$\boxed{\mathcal{L}_{\text{OPO}}=-\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[\omega_{\alpha}(y)\,v(y)\right]+\frac{\mu}{2}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[v(y)^{2}\right]}$ (9)
where $\mu=1/\eta$ is the stiffness parameter, with closed-form solution $v^{*}(y)=\omega_{\alpha}(y)/\mu$.
4.5 Gradient Dynamics: Why OPO Does Not Saturate
Differentiating Eq. (9) w.r.t. $v_{\theta}(y)$:
$\nabla_{v}\mathcal{L}_{\text{OPO}}=-\omega_{\alpha}(y)+\mu\,v_{\theta}(y)$ (10)
with unique equilibrium $v^{*}(y)=\omega_{\alpha}(y)/\mu$. This linear relationship has three desirable properties:
  1. 1.
    Stable Equilibrium: Strict convexity in $v_{\theta}$ yields a unique stationary point.
  2. 2.
    No Saturation: Unlike sigmoid-based losses, the gradient does not vanish as $v$ grows—it increases proportionally to the distance from equilibrium.
  3. 3.
    Orthogonal Control: Changing $\alpha$ (sampling) affects only $\omega_{\alpha}$; changing $\mu$ (optimization) affects only the curvature. The two do not interact.
4.6 Implementation: Log-Ratio Approximation
The ratio $v_{\theta}=\pi_{\theta}/\pi_{\mathrm{ref}}-1$ is not directly available from LLM logits. However, for the log-ratio $\Delta_{\theta}=\log\pi_{\theta}-\log\pi_{\mathrm{ref}}$ (which is directly computable), the first-order Taylor expansion $v=e^{\Delta}-1\approx\Delta$ holds within the trust region. This yields the implementable form:
$\boxed{\mathcal{L}_{\text{OPO}}^{\text{log}}\approx-\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\big[\omega_{\alpha}(y)\,\Delta_{\theta}(y)\big]+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\big[\Delta_{\theta}(y)^{2}\big]}$ (11)
This retains the linear dynamics of Euclidean geometry while operating entirely in log-space. The approximation is justified by on-policy anchoring: setting $\pi_{\mathrm{ref}}=\pi_{\text{old}}$ ensures $\Delta\approx 0$ at the start of each iteration.
4.7 Constructive Derivation: Regression in Ratio Space
We provide an alternative, constructive derivation that arrives at the same OPO loss from pure first principles—no mirror maps required.
Step 1: Define the target offset.
Given the escort weight $\omega_{\alpha}$ and regularization strength $\mu$, the optimal ratio offset is $v^{*}(y):=\omega_{\alpha}(y)/\mu$. This is the ideal deviation from the reference: the quality signal drives the direction; $\mu$ bounds the magnitude.
Step 2: Minimize the distance.
The natural metric in ratio space is mean squared error:
$\mathcal{J}(\theta)=\frac{\mu}{2}\big(v_{\theta}(y)-v^{*}(y)\big)^{2}$ (12)
Step 3: Complete the square.
Expanding and discarding the $\theta$-independent constant $\omega_{\alpha}^{2}/(2\mu)$:
$\mathcal{J}(\theta)=\underbrace{\frac{\mu}{2}\,v_{\theta}^{2}}_{\text{quadratic regularizer}}-\underbrace{\omega_{\alpha}\cdot v_{\theta}}_{\text{linear drive}}+\text{const}\;\;\Longrightarrow\;\;\mathcal{L}(\theta)=-\omega_{\alpha}\,v_{\theta}+\frac{\mu}{2}\,v_{\theta}^{2}$ (13)
which is exactly the OPO objective (Eq. 9). The completing-the-square identity provides a constructive proof: MSE regression in ratio space is gradient-equivalent to the linear-plus-quadratic OPO loss.
Remark 4.5 (Equivalence of Perspectives).
[…]
Remark 4.6 (Analogy with Quantum Amplitudes).
A physical analogy: in quantum mechanics, superpositions are linear in amplitude ($\psi$), while measurements are quadratic in amplitude. Similarly, OPO operates in the ratio space $v$ (“amplitude”), where preference signals superpose linearly, and regularization is quadratic. KL-based methods operate in log-probability (“energy”) space, where strong signals lead to exponential collapse. The amplitude picture explains why OPO sustains non-saturating gradients in high-confidence regimes.
5 Theoretical Analysis
The preceding section derived OPO from two complementary perspectives (mirror descent and ratio regression). We now prove that OPO’s key structural properties—orthogonal control and global contraction—hold rigorously.
5.1 Formal Decoupling of Geometry Axes
All analysis is conducted in the function space of ratio outputs $v\in L^{2}(\pi_{\mathrm{ref}})$, which captures the intrinsic objective geometry. We first bridge to parameter space.
Lemma 5.1 (Parameter Space Gradient Projection).
Let $\mathcal{L}(v_{\theta})$ be a functional of the policy output $v_{\theta}(y)$. The gradient in parameter space $\theta$ is the projection of the function-space gradient $\nabla_{v}\mathcal{L}$ onto the tangent space of the policy manifold:
$\nabla_{\theta}\mathcal{L}(\theta)=\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[\nabla_{v}\mathcal{L}(v_{\theta}(y))\cdot\nabla_{\theta}v_{\theta}(y)\right]$ (14)
Thus, the geometric properties of $\nabla_{v}\mathcal{L}$ directly determine the driving force in parameter space, modulated by the policy’s local sensitivity $\nabla_{\theta}v_{\theta}$.
With this link established, we characterize the geometry of $\mathcal{L}(v)$.
Theorem 5.2 (Decoupled Sampling and Optimization Geometry).
Consider the OPO objective $\mathcal{L}(v)=-\sum_{y}\omega_{\alpha}(y)v(y)+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}[v(y)^{2}]$. In the space of ratio functions $v\in L^{2}(\pi_{\mathrm{ref}})$:
  1. 1.
    The first-order driving force depends on the sampling parameter $\alpha$:
    $\nabla_{v}\mathcal{L}=-\omega_{\alpha}(y)+\mu v(y)$ (15)
  2. 2.
    The second-order curvature is independent of $\alpha$:
    $\nabla^{2}_{v}\mathcal{L}=\mu I$ (16)
Proof.
Direct differentiation yields a constant Hessian $\mu I$, which is independent of $\alpha$. ∎
Remark 5.3 (Scope of Orthogonality).
The structural orthogonality in Theorem 5.2 holds exactly in the ratio function space $L^{2}(\pi_{\mathrm{ref}})$. In parameter space $\theta$, interaction may arise through policy parametrization, since $\nabla_{\theta}v=\nabla_{\theta}(\exp(\Delta)-1)$ depends on the training trajectory influenced by $\alpha$. However, this coupling is a property of the parametrization, not the intrinsic objective curvature: the Hessian $\nabla^{2}_{v}\mathcal{L}=\mu I$ remains structurally independent of $\alpha$ regardless of the parameter mapping.
A natural concern is whether the stop-gradient operator in $\omega_{\alpha}$ (Remark 4.3) introduces bias. The following lemma shows it does not, at the anchor point:
Lemma 5.4 (Exactness of Surrogate Gradient at Anchor Point).
Let $\mathcal{L}_{\mathrm{true}}(\theta)$ be the exact objective without stop-gradient on $\omega_{\alpha}$, and $\mathcal{L}_{\mathrm{OPO}}(\theta)$ the surrogate with $\operatorname{sg}$. Under on-policy anchoring $\pi_{\mathrm{ref}}=\pi_{\mathrm{old}}$, at the start of each iteration ($\theta=\theta_{\mathrm{old}}$):
$\left.\nabla_{\theta}\mathcal{L}_{\mathrm{OPO}}(\theta)\right|_{\theta=\theta_{\mathrm{old}}}=\left.\nabla_{\theta}\mathcal{L}_{\mathrm{true}}(\theta)\right|_{\theta=\theta_{\mathrm{old}}}$ (17)
Furthermore, the gradient error grows only as $\mathcal{O}(\|\pi_{\theta}-\pi_{\mathrm{old}}\|)$.
Proof.
By the product rule, the true function-space gradient contains a cross-term $(\nabla_{v}\omega_{\alpha})\,v$ that the surrogate omits. At the anchor point $\pi_{\theta}=\pi_{\mathrm{old}}$, the ratio offset is exactly $v=\pi_{\theta}/\pi_{\mathrm{old}}-1=0$, so this cross-term vanishes. For $\theta\neq\theta_{\mathrm{old}}$, the error is bounded by $\|(\nabla_{v}\omega_{\alpha})\,v\|=\mathcal{O}(\|v\|)=\mathcal{O}(\|\pi_{\theta}-\pi_{\mathrm{old}}\|)$, which is small within the trust region. ∎
This confirms that stop-gradient is not an engineering hack but a first-order exact approximation: the surrogate and true objectives agree to first order at every on-policy anchor. Under sufficiently small trust-region updates (i.e., $\|v\|$ bounded by the $\chi^{2}$ penalty), the cumulative bias across iterations remains second-order small, since each anchor reset re-zeroes the first-order error term.
Corollary 5.5 (Global Contraction and Linear Convergence).
Under the assumption that the policy class can represent $v^{*}$ and that local linearization of the map $\theta\mapsto v_{\theta}$ holds near the optimum, gradient descent in $v$-space on the OPO objective follows a linear dynamical system:
$v_{k+1}-v^{*}=(1-\eta\mu)(v_{k}-v^{*})$ (18)
where $v^{*}=\omega_{\alpha}/\mu$. For step size $0<\eta<2/\mu$, this system exhibits global contraction to the unique equilibrium. The contraction rate $|1-\eta\mu|$ depends solely on the optimization stiffness $\mu$, confirming that convergence speed is structurally decoupled from the target distribution sharpness $\alpha$.
We use the term “orthogonalized” to emphasize this control decoupling.
Remark 5.6 (Contrast with KL-Based Methods).
In KL-regularized objectives (e.g., DPO), the effective loss locally resembles a logistic function $\mathcal{L}\approx-\log\sigma(\text{margin})$. The local curvature scales as $\beta^{2}\sigma(m)(1-\sigma(m))$, where $m$ is the logit margin. This curvature depends on both the temperature $\beta$ and the current data margin $m$. Thus, changing $\beta$ simultaneously alters both the gradient magnitude and the local stability profile, entangling the two design axes. OPO’s constant curvature $\mu$ avoids this data-dependent instability.
5.2 Comparison with Parametric L2 Regularization
A natural question is whether OPO differs fundamentally from standard Policy Gradient with L2 parameter regularization (weight decay).
$\mathcal{L}_{\text{L2-PG}}=-\mathbb{E}[A\log\pi_{\theta}]+\frac{\lambda}{2}\|\theta-\theta_{0}\|^{2}$ (19)
While both methods induce a quadratic penalty, they operate in fundamentally different spaces:
  1. 1.
    Functional vs. Parametric Prior: OPO’s penalty $\frac{\mu}{2}\mathbb{E}[v^{2}]$ acts in the function space of probability ratios. It penalizes deviations only where the policy assigns probability mass. L2 parameter regularization acts in the weight space. Two different policies can have the same L2 distance, and small weight changes can cause large output deviations (and vice-versa).
  2. 2.
    Adaptive Constraint: OPO’s regularization provides a “linear restoring force” directly on the probability ratio. If the model is confident (large $v$), the $\chi^{2}$ penalty generates a strong, specific gradient signal to pull that specific output back. Parametric L2 is blind to the output; it uniformly decays all weights regardless of their contribution to the current output confidence.
  3. 3.
    Learning Dynamics: Empirically, Functional $\chi^{2}$ regularization allows the model to shift mass aggressively where advantages are high, while maintaining stability. Weight decay suppresses all weight magnitudes, which can inadvertently hinder the learning of sharp, high-confidence reasoning chains required for math tasks.
5.3 Quantitative Gradient Analysis
We provide a quantitative comparison of the gradient signal in high-confidence regimes, proving the non-saturation property.
Gradient Saturation in DPO.
For logistic-based losses (DPO), the gradient magnitude with respect to the margin $m$ is $|\nabla_{m}\mathcal{L}|=|\beta\sigma(m)(1-\sigma(m))|$. Since $\sigma(1-\sigma)\leq 1/4$:
$|\nabla_{\text{DPO}}|\leq\frac{\beta}{4}$ (20)
Crucially, as the margin $m\to\infty$, $|\nabla_{\text{DPO}}|\to 0$ exponentially.
Linear Gradient Dynamics in OPO.
For OPO, the gradient is $\nabla_{v}\mathcal{L}=\mu v-\omega_{\alpha}$. Let $v^{*}=\omega_{\alpha}/\mu$. The gradient magnitude is exactly proportional to the distance from equilibrium: $|\nabla_{v}\mathcal{L}|=\mu|v-v^{*}|$. Therefore, for any non-equilibrium state ($|v-v^{*}|\geq\delta$):
$|\nabla_{\text{OPO}}|\geq\mu\delta$ (21)
This guarantees a linear driving force that does not vanish arbitrarily. The practical consequence is stark: in mathematical reasoning, models frequently assign near-1.0 probability to correct intermediate steps. KL-based methods penalize these high-confidence correct outputs by exponentially shrinking the learning signal, whereas OPO maintains a constant reward-driven pull via its linear dynamics—allowing the model to continue refining even its most confident predictions.
5.4 Algorithm and Flowchart
We provide a visual overview of the OPO process in Figure 1 and the detailed procedure in Algorithm 1.
Current Policy $\pi_{\theta}$ Sampling $(x,y)$ Evaluation $A(y)$ OPO Decomposition 1. Ratio $v=\pi/\pi_{old}-1$ 2. $\alpha$-Weight $\omega_{\alpha}=(e^{A}/\tilde{p})^{1-\alpha}$ 3. Orth. Penalty $\mu\,v^{2}$ Update $\theta$ Next Iteration
Figure 1: Flowchart of the Orthogonalized Policy Optimization (OPO) framework. The “OPO Decomposition” block (green) embodies the two-geometry structure: the $\alpha$-divergence projection (step 2) governs sampling, while the $\chi^{2}$ penalty (step 3) governs optimization. These two axes are structurally orthogonal.
Algorithm 1 Orthogonalized Policy Optimization (OPO)
1:Initial policy $\pi_{\theta}$; parameters $\alpha\in[0,1],\;\mu>0,\;\eta>0$
2:for each iteration do
3:  Anchor: $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ $\triangleright$ On-policy anchoring
4:  Rollout: Sample $(x,y)\sim\pi_{\theta}$; compute rewards $R(x,y)$
5:  Advantage: $A(y)\leftarrow\mathrm{Normalize}(R(x,y))$ $\triangleright$ Group-normalized
6:  Escort Weight: $\omega_{\alpha}(y)\leftarrow\exp\!\big((1{-}\alpha)\,A(y)\big)\cdot\operatorname{sg}\!\big(\pi_{\mathrm{ref}}(y)^{-(1-\alpha)}\big)$ $\triangleright$ Eq. 6
7:  Log-Ratio: $\Delta_{\theta}(y)\leftarrow\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$
8:  Loss: $\mathcal{L}\leftarrow\frac{1}{N}\sum_{y}\big[-\omega_{\alpha}(y)\,\Delta_{\theta}(y)+\frac{\mu}{2}\,\Delta_{\theta}(y)^{2}\big]$ $\triangleright$ $y$: complete response sequence
9:  Update: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$
10:end for
6 Experiments
We conduct an empirical comparison of OPO against three strong baselines—GRPO, GSPO, and DAPO—on mathematical reasoning tasks using the VERL framework.
6.1 Experimental Setup
Model and Data.
We use Qwen3-1.7B as the base model and train on MATH Level 3 problems. All methods use identical training configurations: 4 epochs, batch size 32, learning rate $2\times 10^{-6}$, and 6 rollout generations per prompt.
Baseline Configurations.
  • GRPO (Group Relative Policy Optimization) [4]: Standard token-level policy gradient with group-normalized advantages. Uses vanilla PPO loss mode without a value function critic.
  • GSPO (Group Sentence-level Policy Optimization): A variant of GRPO that applies advantage normalization at the sentence/step level rather than the token level, aiming for more granular credit assignment.
  • DAPO (Dynamic Advantage Policy Optimization): An advanced baseline incorporating four key mechanisms:
    1. 1.
      Clip Higher: Raises the PPO clip upper bound (e.g., $1+\epsilon\to 1.28$) to mitigate entropy collapse and preserve exploration.
    2. 2.
      Dynamic Sampling: dynamically filters "all-correct" or "all-wrong" samples that contribute zero advantage, focusing training on informative samples.
    3. 3.
      Token-Level Policy Gradient: Standardizes advantage weights across tokens in a mini-batch to prevent gradient dilution in long sequences.
    4. 4.
      Overlong Reward Shaping: Applies a length-aware penalty and masks gradients for truncated responses to reduce noise.
  • OPO: The proposed method with $\alpha=0.4$, $\mu=1.0$, on-policy anchoring ($\pi_{\mathrm{ref}}=\pi_{\text{old}}$), and adaptive $\tau$ with range $[0.2,1.5]$. Note: $1-\alpha=0.6$ controls the effective advantage amplification in Eq. (6).
6.2 Results
Overall Performance.
Table 1 summarizes the key metrics across all four algorithms.
Table 1: Comparison of alignment algorithms on Qwen3-1.7B + MATH Level 3. Reward and gradient norm are averaged over the final 20 training steps. Higher gradient norm at convergence, combined with stable (non-exploding) training dynamics, indicates absence of gradient saturation rather than optimization instability.
Algorithm Mean Reward Grad Norm Characteristics
GRPO 0.686 0.61 Standard baseline, stable
GSPO 0.713 0.50 Strong, some gradient decay
DAPO 0.67* 0.22 Conservative, early plateau
OPO (Ours) 0.756 0.90 Best reward, non-saturating gradients
Algorithm Characteristics.
  • GRPO: Serves as the canonical baseline. Achieves reasonable performance (69% acc) with moderate gradient norms. The token-level normalization provides stability but may average out strong signals.
  • GSPO: Improves upon GRPO (71% acc) via sentence-level credit assignment. However, gradient norms (0.50) are lower than OPO’s, suggesting some saturation in high-confidence regimes.
  • DAPO: Exhibits the lowest gradient norms (0.22) and plateaus early. The conservative clipping constraints appear overly restrictive for reasoning tasks where continued learning is beneficial. *Note: Reward fluctuated between 0.59–0.67.
  • OPO: Achieves the highest mean reward (0.756) while maintaining the largest gradient norms (0.90). This empirically validates the theoretical prediction: the $\chi^{2}$ geometry provides non-saturating linear gradients, allowing the model to continue learning where others plateau.
Training Dynamics.
Figure 2 shows the training accuracy (mean reward) over 224 training steps.
  • Early Stage: GRPO exhibits the fastest initial learning, surpassing OPO in the first 50 steps. This is consistent with standard policy gradient methods having aggressive early updates.
  • Late Stage Cross-over: As training progresses (after step 100), GRPO and GSPO begin to saturate. In contrast, OPO maintains a steady improvement rate, eventually overtaking all baselines to reach the highest final accuracy ($\sim$76%). This confirms that the $\chi^{2}$-induced linear drive prevents the “vanishing gradient” problem common in KL-constrained methods as confidence increases.
Figure 2: Training dynamics on Qwen3-1.7B Math RL. Note the cross-over behavior: GRPO (green) starts strong but plateaus, while OPO (red) sustains improvement and achieves the highest final performance.
Gradient Behavior.
Figure 3 illustrates the gradient norm dynamics across training. OPO exhibits consistently higher but stable (non-exploding) gradient norms throughout training, validating the non-saturating property of the $\chi^{2}$ geometry. The stability of these norms rules out the alternative explanation of optimization instability or under-regularization. DAPO shows severely diminished gradients, explaining its early plateau.
Figure 3: Gradient norm comparison. OPO maintains healthy gradient norms (0.9) throughout training, while DAPO (0.22) and GSPO (0.50) exhibit gradient decay.
Entropy Dynamics.
Figure 4 depicts the policy entropy throughout training. OPO maintains consistently higher entropy levels compared to baselines. We attribute this to the granularity of supervision: GRPO and DAPO utilize token-level updates, imposing dense supervision that forces the policy to collapse rapidly into specific phrasings. In contrast, OPO (and GSPO) employs sequence-level updates, where $y$ denotes an entire response trajectory and the loss is averaged over complete sequences rather than individual tokens. This “sparse” supervision aligns the total trajectory probability without micromanaging individual tokens, thereby preserving policy diversity and exploration potential.
Figure 4: Entropy dynamics. OPO maintains higher entropy than baselines, preventing premature mode collapse and enabling sustained exploration throughout the training process.
Summary of Findings.
The experimental results confirm that OPO’s orthogonalized design—combining $\alpha$-weighted sampling with $\chi^{2}$ optimization geometry—prevents gradient saturation and enables continued learning in high-confidence regimes.
7 Discussion
Relation to Existing Methods.
OPO can be viewed as occupying a previously unexplored point in the design space:
  • $\alpha\to 1$ (conservative escort): $\omega_{\alpha}\to 1$, recovering supervised fine-tuning–like uniform weighting.
  • $\alpha\to 0$ (aggressive escort) with KL geometry: the exponential-type weighting resembles KL-regularized preference optimization (e.g., DPO), but inherits gradient saturation.
  • OPO: moderate $\alpha$ (peak-seeking escort) combined with Euclidean ($\chi^{2}$) geometry—non-saturating gradients with advantage-aware sampling.
Computational Cost.
OPO incurs no additional model forward/backward passes compared to DPO/GRPO; the extra bookkeeping for $\omega_{\alpha}$ is lightweight in practice.
Role of the Reference Policy.
In OPO, the reference policy $\pi_{\mathrm{ref}}$ serves solely as an origin defining the coordinate system for the ratio $v_{\theta}=\pi_{\theta}/\pi_{\mathrm{ref}}-1$, rather than as an explicit regularizer via KL divergence. Stability is instead enforced by the quadratic penalty $\frac{\mu}{2}v^{2}$. This decoupling clarifies the distinct roles of anchoring (coordinate system) and regularization (optimization geometry).
Implementation Notes.
The OPO objective (Equation˜9) involves expectations under $\pi_{\mathrm{ref}}$. In practice:
  • For on-policy training where samples come from $\pi_{\mathrm{ref}}=\pi_{\text{old}}$, the batch itself provides an unbiased estimate.
  • For off-policy settings, importance sampling or a separate reference sample set can be used.
Limitations.
The framework introduces hyperparameters ($\alpha$, $\mu$) that may require tuning. Our experiments focus on reasoning tasks; validation on diverse domains (code generation, general instruction following) remains future work. Larger-scale experiments across diverse tasks and model sizes are needed to fully characterize the strengths and limitations of OPO.
8 Conclusion
We have presented Orthogonalized Policy Optimization (OPO), a framework grounded in two independent geometries: an $\alpha$-divergence escort projection that shapes sampling, and a Euclidean mirror map (Pearson $\chi^{2}$) that shapes optimization. The same loss emerges from two complementary derivations—mirror descent and ratio-space regression—confirming it is uniquely determined by the geometry choice. The resulting objective has a constant Hessian, linear non-saturating gradients, and global contraction, addressing a key limitation of KL-based methods in high-confidence reasoning tasks. We believe OPO provides a principled foundation for future work on robust and scalable LLM alignment.
Appendix A Theoretical Proofs and Derivations
A.1 Log-Ratio Approximation Error
Lemma A.1 (Log-Ratio Approximation Error).
Let $\Delta_{\theta}(y)=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$ be the log-ratio, and $v_{\theta}(y)=\exp(\Delta_{\theta}(y))-1$ be the exact ratio deviation. Let $\delta=\|\Delta_{\theta}\|_{\infty}$. For sufficiently small $\delta$ (in particular $\delta<1$, which is enforced by the $\chi^{2}$ trust-region penalty $\frac{\mu}{2}\mathbb{E}[v^{2}]$ bounding $\|\Delta\|_{\infty}$):
$|v_{\theta}-\Delta_{\theta}|\leq\frac{1}{2}\delta^{2}e^{\delta},\quad|v_{\theta}^{2}-\Delta_{\theta}^{2}|\leq\delta^{3}e^{2\delta}$ (22)
Consequently, the difference between the exact ratio loss and the log-approximate loss scales as $O(\mathbb{E}[\Delta_{\theta}^{3}])$.
Proof.
Using Taylor expansion $e^{x}=1+x+\frac{x^{2}}{2}+\dots$, we have $v=\Delta+\frac{\Delta^{2}}{2}+O(\Delta^{3})$. The bounds follow from standard remainder estimation for the exponential function. ∎
A.2 Equivalence to $\chi^{2}$-Constrained Maximization
Proposition A.2 (Lagrange Dual Equivalence).
Consider the problem of maximizing alignment with the target $\omega_{\alpha}$ subject to a functional $\chi^{2}$ trust region:
$\max_{v\in L^{2}(\pi_{\mathrm{ref}})}\mathbb{E}_{\pi_{\mathrm{ref}}}[\omega_{\alpha}(y)v(y)]\quad\text{s.t.}\quad\mathbb{E}_{\pi_{\mathrm{ref}}}[v(y)^{2}]\leq\epsilon$ (23)
The Lagrangian relaxation of this problem is exactly the OPO objective $\mathcal{L}_{\text{OPO}}=-\mathbb{E}[\omega_{\alpha}v]+\frac{\mu}{2}\mathbb{E}[v^{2}]$, where the regularization coefficient $\mu$ is the Lagrange multiplier corresponding to the trust region radius $\epsilon$.
The optimal solution is $v^{*}(y)=\frac{1}{\mu}\omega_{\alpha}(y)$. This duality implies that $\alpha$ purely shapes the objective (alignment direction), while $\mu$ purely enforces the feasibility radius (optimization geometry).
A.3 Proof of Proposition 4.2: $\alpha$-Geometric Interpolation
Proof.
Define the oracle target $P^{*}(y)\propto\exp(A(y))$ and the parameterized family $\rho_{\alpha}(y)\propto\tilde{p}(y)^{\alpha}\,P^{*}(y)^{1-\alpha}$ for $\alpha\in[0,1]$.
Step 1 (Geodesic structure). In the $e$-affine coordinate system of the statistical manifold [11], the log-density of $\rho_{\alpha}$ is:
$\log\rho_{\alpha}(y)=\alpha\log\tilde{p}(y)+(1-\alpha)\log P^{*}(y)-\log Z_{\alpha}$ (24)
where $Z_{\alpha}$ is the normalizing constant. Since $\log\rho_{\alpha}$ is an affine function of $\alpha$ in the natural parameter, $\{\rho_{\alpha}\}_{\alpha\in[0,1]}$ forms an $e$-geodesic (exponential geodesic) connecting $P^{*}$ (at $\alpha=0$) to $\tilde{p}$ (at $\alpha=1$) in the exponential family.
Step 2 (Importance weight). The sampling weight is:
$\omega_{\alpha}(y)=\frac{\rho_{\alpha}(y)}{\tilde{p}(y)}\propto\left(\frac{P^{*}(y)}{\tilde{p}(y)}\right)^{1-\alpha}$ (25)
In the on-policy setting where $P^{*}(y)=\exp(A(y))$ and $\tilde{p}=\pi_{\mathrm{old}}$, this yields $\omega_{\alpha}\propto(\exp(A)/\tilde{p})^{1-\alpha}$, which simplifies to $\exp((1-\alpha)A)$ when the $\tilde{p}$ correction is absorbed by the sampling measure.
Step 3 (Geometric consistency). This interpolation corresponds to the exponential geodesic in the statistical manifold and is consistent with Rényi $\alpha$-divergence interpolation geometry [11]. Under the dually flat structure of exponential families, it coincides with the $e$-projection onto the geodesic submanifold connecting $\tilde{p}$ and $P^{*}$. ∎
A.4 Distributional Stability Guarantees
Proposition A.3 ($\chi^{2}$ Controls Total Variation).
Let $TV(\pi,\pi_{\mathrm{ref}})=\frac{1}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}[|t(y)-1|]=\frac{1}{2}\mathbb{E}[|v(y)|]$. By Jensen’s inequality:
$TV(\pi,\pi_{\mathrm{ref}})=\frac{1}{2}\mathbb{E}[|v|]\leq\frac{1}{2}\sqrt{\mathbb{E}[v^{2}]}$ (26)
Thus, bounding the $\chi^{2}$ norm $\mathbb{E}[v^{2}]\leq\epsilon$ guarantees that the policy remains within a $\sqrt{\epsilon}/2$-radius of the reference in Total Variation distance.